OpenStudy (anonymous):

Tell me if im doing this right 3/5 of 15= 3/5*15 divide 15 into 5 pieces will i get 9

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

i need a little help understanding equivalent chains

OpenStudy (anonymous):

this is the the question.. 2/6=3/9=?=?=10/?=?=?=?/48=?=?